Find a recurrence relation for the number of bit strings of length n that contain a pair of consecutive 0s. Find also the initial conditions?
1
Expert's answer
2020-06-02T17:32:34-0400

Let "a_n" represents the number of bit strings of length n that contain a pair of consecutive 0's.

a) First Case: The bit string is a sequence ending in 1, then the bit string of length n-1 (ignore the last 1) has "a_{n-1}" possible strings that contain a pair of consecutive 0's.


Second case: The bit string is a sequence ending in 10, then the bit string of length n-2 (ignore the last two digit 10) has "a_{n-2}" possible strings that contain a pair of consecutive 0's.


Third case: The bit string is a sequence ending in 00. There are "2^{n-2}" bit strings of length n-2 and thus there are "2^{n-2}" bit of string n-2 followed by 00.


Adding the number of sequences of all three cases, we then obtain:

"a_n = a_{n-1}+a_{n-2}+ 2^{n-2}" .


b) When "n=0", there are exactly 0 bit strings with a pair consecutive 0's (since the empty string does not contain a pair of consecutive zeros).

"a_0 = 0".

When "n=1", there are exactly 0 bit strings with a pair of consecutive 0's (since the string contains only 1 element).

So, "a_1 = 0".

When "n=2" , there is exactly 1 bit string with a pair consecutive 0's (namely the bit string 00). When we use the recurrence relation in part(a), we note that "a_2 = a_1+a_0 +2^{2-2} = 0 + 0 +1 = 1". Thus the recurrence relation holds for "n=2" and thus "n=2" does not need to be an initial condition.